Voice assimilation in Hungarian Péter Szigetvári, Eötvös Loránd University, szigetva@isis.elte.hu

(1) The data monomorphemically

afgán  [vg] ‘Afghan(istani)’  Agfa  [kf] id.
Macbeth [gb] id.  Leeds  [ts] id.
Stuttgart [dg] id.  vodka  [tk] id.

across a morpheme boundary

háztól [st] ‘house-abl’ hatból [db] ‘six-elat’
polgazdként [stk] ‘as political economy’ lisztből [zdb] ‘flour-elat’
smaragdféle [ktf] ‘emerald like’  receptbe [bdb] ‘recipe-illat’

across a word boundary

kék bélyeg [gb] ‘blue stamp’  zöld kutya [ltk] ‘green dog’

(2) The representation of voicing

a. laryngeal elements: L and H. L=[slack vocal folds], H=“[stiff vocal folds]” (KLV 1990 : 216), but (i) these two features cannot cooccur in a segment (Halle & Stevens 1971, Kaye 1995 : 325), i.e. they are not really unary features, (ii) if H is responsible for aspiration (Harris 1994 : 133ff), it is not [stiff vocal folds] but [spread glottis]

b. a binary laryngeal contrast is either H vs. ∅ or ∅ vs. L (cf. Harris 1994 :

133ff). Hungarian (probably) has ∅ vs. L c. regressive VA is the spreading of L from the trigger ([tb] → [db]) or the

delinking of L from the target ([zt] → [st]) d. The meaning of L

L does not mean that the segment containing it exhibits vocal fold vibration, but that it is markedly voiced: an effort has to be made to produce vocal fold vibration (Chomsky & Halle 1968 : 300f, Hayes 1984). Thus nonobstuents do not contain L, they exhibit spontaneous voicing, which is phonologically irrelevant.

e. problems:

α. why does L delink? more on this in (3) β. what stops the spreading of L?

standard assumption: L can only spread if the target has a Laryngeal node, L links to the skeleton by the mediation of this node. f. If (2d) is true, all and only obstruents have Lar, i.e. either [−son] or

Lar is redundant in the representation of obstruents. Proposal: Lar is equivalent to h, the [noise] element inherent in obstruents (Harris 1990 : 263). Representations:

voiced obstruent  voiceless obstruent  sonorant
x  x  x
h  h

L

g. L spreads backwards until it finds an h on the superficially adjacent segment. (In Russian h-less nonvowels can be skipped: o[d] mzdy ‘from the bribe’ (Hayes 1984 : 320); vowels cannot: taza *[daza] ‘of the pelvis’.)

h. Voiceless sonorants are aspirated (Lombardi 1995b : 51), from which it follows that H is not dominated by Lar (h here). Note that regressive spreading of aspiration is not as obvious as that of voicing (Greek spelling has ϕθ and χθ, but πϕ, τθ, κχ).

(3) The site a. becoming voiceless = loss of L b. lenition site:

L is unlicensed in coda (Lombardi 1995a) L is unlicensed in coda and pre-empty-nuclear onset (Brockhaus 1995) ablak ‘window’, abrak ‘fodder’ vs. paplan ‘quilt’, apró ‘tiny’; hab ‘foam’, habnak ‘foam-dat’ vs. ha[p]tól ‘foam-abl’; ha[�]ma ‘onion’ vs. fi[c]ma ‘prepuce’; bab ‘bean’, bableves ‘bean soup’ vs. babszem [ps] ‘a piece of bean’ ⇒ loss of L only occurs before a voiceless obstruent:

*x  x^{h  h} L

c. problem: the constraint above makes an implicit reference to the absence of a feature: the first position cannot license L only if the second has h without L.

novi sad workshop on gp, 6–8 december 1997 3

d. conclusions:

α. the site of VA cannot be captured by making reference to syllabic

constituents: neither coda, nor pre-empty-nuclear onset position β. in the formulation of VA the absence of L must be referred to

e. a temporary solution: of two adjacent h’s the first is identified with the second (“identified with its governor” (Brockhaus 1995 : 133; though whether it is its governor is doubtful in most cases, cf. Lukács 1997)

(4) Nonobstruents in VA: v, j, h a. prevocalically v [�], j and h are sonorants b. non-prevocalically v and h are obstruents [v] and [x], j is an obstruent

[�]/[ç] only when not next to a vowel c. data

ásvány [��] ‘mineral’ elvtárs [lft] ‘comerade’
kedvtelen [tft] ‘depressed’ két Wrangler [tvr] ‘two pairs of W. jeans’
céhbeli [xb] ‘guildsman’ egyház [ch]  ‘church’
bojt  [jt] ‘tassel’  rakj át [kj]  ‘put-imp over’
dobj  [b�] ‘throw-imp’ rakj [kç]  ‘put-imp’
dobj ki [pçk] ‘throw out’ rakj be [g�b]  ‘put in’
szomj [m�] ‘thirst’ férj [r�]/[rç] ‘husband’
szomjról [m(�)r]‘thirst-delat’ férjről [r(�)r]/*[rçr]‘husband-delat’

d. problems:

α. how do v and h become obstruents in non-prevocalic position? β. why does [x] not become voiced (if an obstruent)? γ. why is obstruentized [v] and [�] voiced (where does their L come

from)? e. attempts at answers and further details:

ad α. does h get into the representation? if yes, what is its source and

is this a case of coda-fortition? if no, how can [v] become voiceless ad β. assuming that [h] is H (and not h, cf. Szigetvári 1996), it does

not get voiced because H and L are incompatible in Hungarian (i.e. there are no voiced aspirates; Szigetvári 1997)

ad γ. ??? (cf. Cyran 1997 : 198ff; but he describes a historical fortition with reanalysis, while here we have a synchronic alternation)

4 szigetvári, va in hungarian

(5) Claims made a. there is no Lar node, its job is taken over by h b. H is not [stiff vocal folds], but [spread glottis] c. H is not dominated by h (i.e. by Lar) d. [h] is not h, but H (h is [s], especially if there is no R) e. VA (in Hungarian) is not describable with reference to syllabic

constituents, a linear formulation is simpler and more precise (cf. Rubach 1996)

references

Brockhaus, Wiebke. 1995. Final devoicing in the phonology of German. Linguistische

Arbeiten 336. Tübingen: Niemeyer. Chomsky, Noam and Morris Halle. 1968. The sound pattern of English. New York: Harper

& Row. Cyran, Eugeniusz. 1997. Resonance elements in phonology: A study in Muster Irish. Lublin:

Folium. Halle, Morris and Kenneth Stevens. 1971. A note on laryngeal features. Quarterly Progress

Report 101 : 198–212. Cambridge, Mass.: Research Laboratory of Electronics, MIT. Harris, John. 1990. Segmental complexity and phonological government. Phonology 7 :

255–300. Harris, John. 1994. English sound structure. Oxford: Blackwell. Hayes, Bruce. 1984. The phonetics and phonology of Russian voicing assimilation. In Mark

Aronoff and Richard T. Oehrle (eds.) Language sound structure. Studies in phonology presented to Morris Halle by his teacher and students. Cambridge, Mass.: The MIT Press. 318–328.

Kaye, Jonathan. 1995. Derivations and interfaces. In Jacuqes Durand and Francis Katamba.

Frontiers of phonology: Atoms, structures, derivations. Harlow: Longman. 289–332. Kaye, Jonathan, Jean Lowenstamm and Jean-Roger Vergnaud. 1990. Constituent structure

and government in phonology. Phonology 7 : 193–231. Lombardi, Linda. 1995a. Laryngeal neutralization and syllable wellformedness. NLLT 13 :

39–74. Lombardi, Linda. 1995b. Laryngeal features and privativity. The Linguistic Review 12 :

35–59. Lukács, Ágnes. 1997. Interonset govenment and Hungarian: some theoretical considerations.

The Odd Yearbook 1997 : 137–149. Rubach, Jerzy. 1996. Nonsyllabic analysis of voice assimilation in Polish. Linguistic Inquiry

27 : 69–110. Szigetvári, Péter. 1996. Laryngeal contrasts and problematic representations. The Even

Yearbook 2 : 97–110. Szigetvári, Péter. 1997. Why [h] does not get voiced. Ms., Eötvös Loránd University.